The length of a rectangle is twice its width. If the area of the rectangle is 98 cm2., find it's perimeter.
Answer by ewatrrr(24785)

 
Hi,
.
2L + 2w = 98cm^2
L + w = 49 cm^2
(2w) + w = 49
3w = 49
w = 16 1/3 cm and L = 32 2/3 cm